In the laboratory a coffee cup calorimeter, or constant pressure calorimeter, is frequently used to determine the specific
Answer #1

Solution :-

Part 1) mass of water =83.17 g

Mass of silver = 63.04 g

Temperature change of water = 25.59C – 22.31 C = 3.28 C

Temperature change of silver = 98.96 C – 25.59 C = 73.37 C

Specific heat capacity of calorimeter = 1.61 J/C

Step 1) Calculating the heat absorbed by the calorimeter and water

q (water + cal) = (m*c*delta T) + (c cal*Delta T)

                           =(83.17 g * 4.184 J/gC * 3.28 C)+(1.61 J/C * 3.28 C)

                           =1147 J

Heat given off by silver is same as heat absorbed by calorimeter and water

q cal = 1147 J

calculating the specific heat of silver

q= m*c*delta T

c= q/ (m*delta T)

   = 1147 J / (63.04 g * 73.37 C)

   =0.248 J/gC

Therefore the specific heat of silver is 0.248 J/goC

Part 2) mass of water =77.16 g

Temperature change of water = 32.41 C- 23.52 C =8.89 C

Mass of nickel =98.72 g

Temperature change of nickel = 98.18 C – 32.41 C = 65.77 C

Calculating the heat absorbed by water

q= m*c*delta T

= 77.16 g * 4.184 J/gC * 8.89 C

   = 2870 J

Calculating heat lost by nickel

q= m*c*delta T

=98.72 g * 0.445 J/gC * 65.77 C

=2890 J

Calculating heat absorbed by calorimeter

q cal = q nickel – q water

          =2890 J – 2870 J

          =20.0 J

Calculating specific heat of calorimeter

c cal = q cal / Delta T

          = 20.0 J / (8.89 C)

          =2.25 J/C

Therefore the calorimeter constant is 2.25 J/oC